# Acceleration-induced transport of quantum vortices in joined atomtronic circuits

### v1

@@ Abstract
Persistent currents–inviscid quantized flow around an atomic circuit–are a crucial building block of atomtronic devices. We investigate how acceleration influences the transfer of persistent currents between two density-connected, ring-shaped atomic Bose-Einstein condensates, joined by a tunable weak link that controls system topology. We find that the acceleration of this system modifies both the density and phase dynamics between the rings, leading to a bias in the periodic vortex oscillations studied in T. Bland et al., Phys. Rev. Research 4 4, 043171 (2022). Accounting for dissipation suppressing such vortex oscillations, the acceleration facilitates a unilateral vortex transfer to the leading ring. We analyze how this transfer depends on the weak-link amplitude, the initial persistent current configuration, and the acceleration strength and direction. Characterization of the sensitivity to these parameters paves the way for a new platform for acceleration measurements.
@@ 1 Introduction
Unlike quantum gases in conventional harmonic traps, ring-traps can support multiply quantized vortices known as persistent currents [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] (for a recent review see Ref. [20]), superfluid flow that can circulate indefinitely without dissipation. The ability to engineer and control such currents through external perturbations like weak links and potential barriers is critical for creating robust atomtronic devices, which are the quantum analogs of electronic circuits. While numerous proposals suggest using persistent currents for precision rotation measurements [21, 22, 23], existing ultracold atomic accelerometers and gravimeters [24, 25, 26, 27, 28, 29, 30, 31] do not yet utilize persistent currents as their core mechanism.
@@ 2.1 Double-ring geometry and numerical model
Here, $g_{\text{2D}}=\sqrt{2\pi}\hbar^{2}a_{s}/Ml_{z}$ is the effective two-dimensional two-body local interaction coupling strength, $a_{s}$ is the background $s$-wave scattering length, $M$ is the atomic mass, and $l_{z}=\sqrt{\hbar/M\omega_{z}}$ is the harmonic oscillator length in $z$. As an example realistic system, we consider parameters inspired from the experiment of Ref. [34], but with atom number $N=10^{6}$. This corresponds to 23Na atoms with $a_{s}=2.75$ nm, and we also choose $\omega_{r}=2\pi\times 134$ Hz, $R=22.6~{}\mu$m, and a potential width of $l_{b}=3.45~{}\mu$m. This gives a chemical potential of $\mu_{\text{2D}}=2000\,$Hz. The density is normalized to the total atom number $\iint\text{d}x\text{d}y\,|\psi(x,y,t)|^{2}=N$.
To solve Eq. (4), we employ a split-step Fourier method with grid resolution $N_{x}=N_{y}=256$, a spatial size $L_{x}=L_{y}=100l_{z}$, and a time step $\Delta t=2.5\times 10^{-6}\omega_{z}^{-1}$. We have verified that our results are unchanged after doubling the number of grid points in each direction.
Figure 1: Double-ring system with a persistent current under constant horizontal acceleration $a=0.01g$. (a) Density $n=|\psi|^{2}$ and (b) phase distributions taken at time $t=0$. (c) Density and (d) phase distributions relative to the static condensate ($a=0$). The red cross denotes the antivortex position, and black arrows denote the flow direction induced by the vortex. In (d), Green arrows denote the additional flow induced by acceleration seen in the leading ring.
@@ 2.2 Accelerating double-ring system with a closed barrier
To understand the role of acceleration on our system, we first investigate the real-time dynamics in a double-ring system whilst keeping the barrier closed ($V_{0}=0$) with an antivortex imprinted into the leading ring, with respect to the acceleration direction. A numerical demonstration of the role of acceleration is shown in Fig. 1, for a specific example of $a=a_{x}=0.01\,g$ (relative to Earth’s gravity $g=9.81$ m/s2), highlighting that the key modification compared to the previously studied static case is a density redistribution. The general behavior is quite clear from the Thomas-Fermi solution to Eq. (4), where the density is lower in the forwardmost part of the system. This is reminiscent of the classical situation, in how a water surface inclines in a water tank moving with constant acceleration.
In the presence of a persistent current, the density variation also leads to a phase variation away from the usual azimuthal $2\pi$ winding, as shown in Fig. 1(d), where the phase is plotted relative to the static case. This additional phase variation appears in order to satisfy the hydrodynamic continuity equation. Hence, the flow velocity is greater in a less dense region–the forwardmost part of the system–and here we have faster phase accumulation. This build-up of phase at the bottom of the leading ring resembles a Magnus-like effect [35], and if we instead imprint a vortex with positive circulation, this effect flips about the $y$-axis. This phase effect also occurs for a single ring under acceleration.

### v2

@@ Abstract
Persistent currents–inviscid quantized flow around an atomic circuit–are a crucial building block of atomtronic devices. We investigate how acceleration influences the transfer of persistent currents between two density-connected, ring-shaped atomic Bose-Einstein condensates, joined by a tunable weak link that controls system topology. We find that the acceleration of this system modifies both the density and phase dynamics between the rings, leading to a bias in the periodic vortex oscillations studied in T. Bland et al., Phys. Rev. Research 4 4, 043171 (2022). Accounting for dissipation suppressing such vortex oscillations, the acceleration facilitates a unilateral vortex transfer to the leading ring. We analyze how this transfer depends on the weak-link amplitude, the initial persistent current configuration, and the acceleration strength and direction. Characterization of the sensitivity to these parameters paves the way for a new platform for acceleration measurements, for which we outline a proof-of-concept ultracold double-ring accelerometer.
@@ 1 Introduction
Unlike quantum gases in conventional harmonic traps, ring-traps can support multiply quantized vortices known as persistent currents [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] (for a recent review see Ref. [20]), superfluid flow that can circulate indefinitely without dissipation. The ability to engineer and control such currents through external perturbations like weak links and potential barriers is critical for creating robust atomtronic devices, which are the quantum analogs of electronic circuits. While numerous proposals suggest using persistent currents for precision rotation measurements [21, 18, 22] and gyroscopy [23], existing ultracold atomic accelerometers and gravimeters [24, 25, 26, 27, 28, 29, 30, 31] do not yet utilize persistent currents as their core mechanism.
@@ 2.1 Double-ring geometry and numerical model
Here, $g_{\text{2D}}=\sqrt{2\pi}\hbar^{2}a_{s}/Ml_{z}$ is the effective two-dimensional two-body local interaction coupling strength, $a_{s}$ is the background $s$-wave scattering length, $M$ is the atomic mass, and $l_{z}=\sqrt{\hbar/M\omega_{z}}$ is the harmonic oscillator length in $z$. As an example realistic system, we consider parameters inspired from the experiment of Ref. [34], but with atom number $N=10^{6}$. This corresponds to 23Na atoms with $a_{s}=2.75$ nm, and we also choose $\omega_{r}=2\pi\times 134$ Hz, $\omega_{z}=2\pi\times 550$ Hz, $R=22.6~{}\mu$m, and a potential width of $l_{b}=3.45~{}\mu$m. This gives a chemical potential of $\mu_{\text{2D}}=2000\,$Hz. The density is normalized to the total atom number $\iint\text{d}x\text{d}y\,|\psi(x,y,t)|^{2}=N$.
To solve Eq. (4), we employ a split-step Fourier method with grid resolution $N_{x}=N_{y}=256$, a spatial size $L_{x}=L_{y}=100l_{r}$, and a time step $\Delta t=2.5\times 10^{-6}\omega_{r}^{-1}$, where $l_{r}=\sqrt{\hbar/M\omega_{r}}$ is the radial harmonic oscillator length. We have verified that our results are unchanged after doubling the number of grid points in each direction.
Figure 1: Effect of acceleration on a double-ring system with a persistent current. (a) Density $n=|\psi|^{2}$ and (b) phase $\Phi$ distributions at zero acceleration. (c) and (d) show the density and phase under horizontal acceleration $a=0.01g$ in the positive $x$-direction. To visualize the induced changes, (e) and (f) present the relative differences between the columns. The red cross marks the antivortex position, and black arrows indicate the flow direction induced by the vortex. In (f), green arrows show the additional flow induced by acceleration in the leading ring.
@@ 2.2 Accelerating double-ring system with a closed barrier
To understand the role of acceleration on our system, we first investigate the effect of acceleration in a double-ring system while keeping the barrier closed ($V_{0}=0$) with an antivortex imprinted into the leading ring, with respect to the acceleration direction. A numerical demonstration of the role of acceleration is shown in the second column of Fig. 1, for a specific example of $a=a_{x}=0.01\,g$ (relative to Earth’s gravity, $g=9.81$ m/s2), taken at $t=0$, after solving Eq. (4) in imaginary time. This highlights that a key modification compared to the previously studied static case, shown in the first column, is a density redistribution. The general behavior is quite clear from the Thomas-Fermi solution to Eq. (4), where the density is lower in the forwardmost part of the system. This is reminiscent of the classical situation, in how a water surface inclines in a water tank moving with constant acceleration.
In the presence of a persistent current, the density variation also leads to a phase variation away from the usual azimuthal $2\pi$ winding, as shown in Fig. 1(f), where the phase is plotted relative to the static case. This additional phase variation appears in order to satisfy the hydrodynamic continuity equation. Hence, the flow velocity is greater in a less dense region–the forwardmost part of the system–and here we have faster phase accumulation. This build-up of phase at the bottom of the leading ring resembles a Magnus-like effect [35], and if we instead imprint a vortex with positive circulation, this effect flips about the $y$-axis. This phase effect also occurs for a single ring under acceleration.
@@ 2.3 Asymmetric undamped persistent current oscillations
In Fig. 2, the time at which the barrier ramp-down sequence is initiated affects the final position of the vortex. In the static case, it is known that the oscillations halt as soon as $V_{0}$ becomes smaller than $\mu_{\text{2D}}$. As we will discuss later in the damped regime [Fig. 4], the critical barrier amplitude required to suppress oscillations may be even lower under acceleration. If the goal is to control the final state of the vortex, then this can be achieved by calibrating the maximum barrier amplitude to be as close to the chemical potential as possible. Furthermore, since the oscillation period is independent of acceleration, the optimal barrier closing time can be predetermined.